To what Celsius temperature must 63.5 mL of krypton gas at 10.0°C be changed so the volume will triple? Assume the pressure and the amount of gas are held constant. Enter your answer to four significant figures?
Answer #1

PV = nRT

P , n and R are constant

V1 = 63.5 mL

T1 = 273 + 0 = 283 K

V2 = 3V1

T2 = ?

V1 / T1 = V2 / T2

63.5 / 283 = 190.5 / T2

0.224 = 190.5 / T2

T2 = 190.5 / 0.224 = 850.4 K

Temperature in degree celcius = 850.4 - 273 = 577.40 C